# Disentanglement in Implicit Causal Models via Switch Variable

### v1

@@ Abstract
Learning causal representations from observational and interventional data in the absence of known ground-truth graph structures necessitates implicit latent causal representation learning. Implicitly learning causal mechanisms typically involves two categories of interventional data: hard and soft interventions. In real-world scenarios, soft interventions are often more realistic than hard interventions, as the latter require fully controlled environments. Unlike hard interventions, which directly force changes in a causal variable, soft interventions exert influence indirectly by affecting the causal mechanism. In this paper, we tackle implicit latent causal representation learning in a Variational Autoencoder (VAE) framework through soft interventions. Our approach models soft interventions effects by employing a causal mechanism switch variable designed to toggle between different causal mechanisms. In our experiments, we consistently observe improved learning of identifiable, causal representations, compared to baseline approaches.
Keywords Causality  $\cdot$ Disentanglement  $\cdot$ Soft Intervention  $\cdot$ Variational Inference
@@ 1 Introduction
Causal representation learning provides the tantalizing possibility of elevating conventional deep learning models from black boxes to gray boxes such that we can obtain robust, transferable, and explainable representations Schölkopf (2019). In contrast, in statistical representation learning, the representations do not generally have a meaningful structure. Although we can represent statistical models with a Directed Acyclic Graph (DAG), just as in causal models, the connections in such a graph are volatile and cannot be generalized to different domains. Furthermore, in causal models we can perform interventions and counterfactuals on the variables which can help the model to provide meaningful responses for situations that it has never seen before.
Causal models can be expressed by both graphical models and structural causal models (SCMs) Pearl (2005). A causal graphical model is a DAG in which there is an edge between variables such as $X\rightarrow Y$ if $X$ is a parent of $Y$. We call a variable endogenous if it has a parent in the graph and exogenous if it has no parents. Exogenous variables usually correspond to uncertainties and noise in the data Pearl et al. (2016). In graphical causal models we can only make inferences about variables which are causally related to each other, but we cannot make inferences about how they are related to each other. SCMs are a way to describe this relationship using functions.
One of the long-standing challenges in the field is how to recover the ground-truth causal graph of a system from observations alone. This is known as the identifiablity of causal models problem. Without identifiablity, we cannot claim the learnt representations are causal, since statistical models can also be represented by DAGs where the edges are not necessarily causal. To see this, let’s consider three random variables $X,Y,Z$ whose joint distribution is factorized as $p(X,Y,Z)=p(Y|X)p(Z|X)p(X)$. According to the faithfulness assumption, the conditional independencies in the joint distribution of variables must also be entailed in its corresponding graphYu et al. (2020). However, for this joint distribution, the conditional independence $Z\perp\!\!\!\perp Y|X$ alone can be represented by three different graphs: $Z\rightarrow X\rightarrow Y$, $Y\rightarrow X\rightarrow Z$, and $Y\leftarrow X\rightarrow Z$. With respect to this independence relation, these three graphical models belong to the same Markov Equivalence Class (MEC). Thus, the Markov condition in graphs is insufficient for identifying causal models Schölkopf (2019) and without further assumptions or data, we can only learn a MEC of the causal model. Existing works have made different assumptions about availability of ground-truth causal variables labels Yang et al. (2021a), model parameters Ahuja et al. (2023), availability of paired interventional data Brehmer et al. (2022), and availability of intervention targets Lippe et al. (2022a) to ensure identifiability of causal models.
@@
In a Variational AutoEncoder (VAE) framework, there are generally two approaches for causal representation learning: Explicit Latent Causal Models (ELCMs) Yang et al. (2021a); Ahuja et al. (2023); Yang et al. (2021b); Yu et al. (2019); Lippe et al. (2022a); Lachapelle et al. (2022) and Implicit Latent Causal Models (ILCMs) Brehmer et al. (2022). In ELCMs, the latents are the causal variables and the adjacency matrix of the causal graph is parameterized and integrated in the prior of the latents such that the prior of latents is factorized according to the Causal Markov Condition Schölkopf et al. (2021). This approach in causal representation learning is highly susceptible to being stuck in local minima as it is hard to learn representations without knowing the graph, and it is hard to learn the graph without knowing the representations. ILCMs Brehmer et al. (2022) were introduced to circumvent this “chicken-and-egg” problem by using solution functions, which can implicitly model edges in the causal graph rather than explicitly modeling the entire adjacency matrix of the causal model. In ILCMs the latents are the exogenous variables and the there is no explicit parameterization for the graph.
Let’s consider a simple SCM to understand the motivation behind solution functions:
$\mathcal{Z}_{1}=f_{1}(\mathcal{E}_{1})\quad\&$ $\mathcal{Z}_{2}=f_{2}(\mathcal{Z}_{1},\mathcal{E}_{2}),$ (1)
where $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ are exogenous variables, $f_{1}$ and $f_{2}$ are the actual causal mechanisms, and $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ are the causal variables being modelled. This example SCM is a general case in which the causal variables are an arbitrary nonlinear function of parents and exogenous variables. We can now easily express all the causal variables as a function of exogenous variables:
$\left\{\begin{array}[]{ll}\mathcal{Z}_{1}=f_{1}(\mathcal{E}_{1})=s_{1}(\mathcal{E}_{1},\mathcal{E}_{2})\\ \mathcal{Z}_{2}=f_{2}(f_{1}(\mathcal{E}_{1}),\mathcal{E}_{2})=s_{2}(\mathcal{E}_{1},\mathcal{E}_{2}),\end{array}\right.$ (4)
where $s_{1}$ and $s_{2}$ are the solution functions. We see that $s_{2}$ takes the two exogenous variables as input, as it must in order to map them to the appropriate causal variables. But note that $s_{1}$ also takes this same form, even though in this case, $\mathcal{E}_{2}$ won’t be needed for modeling $\mathcal{Z}_{1}$. We leave it to the solution function to learn which of the exogenous variables are ancestral. Each solution function will thus take in all exogenous variables and learn to ignore some of them. This makes solution functions simpler to define than a full adjacency matrix or even multiple predefined functions, and no knowledge of the graph structure is needed a priori. One might wonder why exogenous variables are chosen to be the latents in ILCMs. The reason is that exogenous variables must be mutually independent in a SCM, hence, we can conveniently factorize the priors of the exogenous variables as $p(\mathcal{E})=\Pi_{i}p(\mathcal{E}_{i})$. The observed variables are obtained as mixture function $g$ of causal variables $\mathcal{Z}$, i.e., $\mathcal{X}=g(\mathcal{Z})$.
In implicit causal representation learning, the task involves recovering the exogenous variables $\mathcal{E}$ from observed variables $\mathcal{X}$ and learning solution functions. In Brehmer et al. (2022), interventions are assumed to be hard, but this is unrealistic and does not align with real-world problems Correa and Bareinboim (2020). Hard interventions involve severing the connections between intervened variables and their parents Pearl et al. (2016). While in soft interventions, the causal variable is still affected by its parents, thus intervention can influence the conditional distribution $p(\mathcal{Z}_{i}|\mathcal{Z}_{pa})$ Correa and Bareinboim (2020). Soft interventions include hard interventions within them as a special case by using the intervention $do(\mathcal{Z}_{i}=z_{i})$.
@@
In this paper, we propose a novel approach for implicit causal representation learning with soft interventions. We will introduce the causal mechanism switch variable as a way of modeling the effect of soft interventions and identifying the causal variables. Our experiments on both synthetic and large real-world datasets, highlight the efficacy of proposed method in identifying causal variables and promising future directions in implicit causal representation learning.
Our key contributions can be summarized as follows:
  1. 1.
@@
    A novel approach for implicit causal representation learning with soft interventions.
  2. 2.
@@
    Employing causal mechanisms switch variable to model the effect of soft interventions.
  3. 3.
    Theory for identifiability up to reparameterization and causal disentanglement from soft interventions.
@@ 2 Related Work
2.1 Causal Representation Learning
Causal representation learning has recently garnered significant attention Schölkopf et al. (2021); Kaddour et al. (2022). The primary challenge in this problem lies in achieving identifiability beyond the Markov equivalence class Schölkopf (2019). Solely relying on observational data necessitates additional assumptions regarding causal mechanisms, decoders, latent structure, and the availability of interventional data Lu et al. (2022); Shen et al. (2022); Yu et al. (2020); Perry et al. (2022); Lachapelle et al. (2022); Ahuja et al. (2023); Zheng et al. (2018); Jaber et al. (2020); Yang et al. (2021a). Recent works have focused on identifying causal models from collected interventional data instead of making strong assumptions about functions of the causal model. Interventional data facilitates identifiability based on relatively weak assumptions Ahuja et al. (2023); Cooper and Yoo (2013); Brehmer et al. (2022); Zhang et al. (2023a); Wendong et al. (2023). This type of data can be further categorized based on whether it involves soft or hard interventions, and whether the manipulated variables are observed and specified or latent. Our focus in this paper is on examining soft interventions, encompassing both observed and unobserved variables.
@@ 2.1 Causal Representation Learning
CausalDiscrepancy Zhang et al. (2023b) Nonlinear Full row rank polynomial Soft Explicit Permutation and Affine
CauCA Wendong et al. (2023) Nonlinear Diffeomorphism Soft Explicit Different based on assumptions
Linear-CD Squires et al. (2023) Linear Linear Hard Explicit Permutation
Scale-I Varici et al. (2023) Nonlinear Linear Hard/Soft Explicit Scale/Mixed
ILCM Brehmer et al. (2022) Nonlinear Diffeomorphism Hard Implicit Permutation and reparameterization
dVAE Locatello et al. (2020) Nonlinear Diffeomorphism Hard Implicit Permutation and reparameterization
SoftCD (ours) Nonlinear Diffeomorphism Soft Implicit Reparameterization
@@
Table 1 presents a comparison of the assumptions and identifiability results between our proposed theory and other related works on causal representation learning with interventions. In causal representation learning with interventions, one approach assumes a given causal graph and concentrates on identifying causal mechanisms and mixing functions. For instance, Causal Component Analysis (CauCA) Wendong et al. (2023) explores soft interventions with a known graph. Alternatively, when the graph is not provided, explicit models seek to reconstruct it from interventional data Cooper and Yoo (2013); Lippe et al. (2022a), potentially resulting in a chicken-and-egg problem in causal representation learning Brehmer et al. (2022). Current methods face the challenge of simultaneously learning the causal graph and other network parameters, especially in the absence of information about causal variables or the graph. Addressing these challenges, Brehmer et al. (2022) recently introduced ILCM, which performs implicit causal representation learning exclusively using hard intervention data. In contrast, our approach introduces a novel method for learning an implicit model from soft interventions. Brehmer et al. (2022) describes methods for extracting a causal graph from a learned implicit model, which could be applied to our method as well. In our experiments, we will compare our method with ILCM and dvae Locatello et al. (2020), given their implicit nature and similar experimental settings and assumptions. Additionally, to showcase the superiority of our method over explicit models, we will employ explicit causal model discovery methods like ENCO Lippe et al. (2022b) and DDS Charpentier et al. (2022), in conjunction with various variants of $\beta$-VAE.
@@ 2.3 Hard interventions vs Soft interventions
The identification of explicit causal models from hard interventions has been extensively explored. Squires et al. (2023) investigate causal disentanglement in linear causal models with linear mixing functions under hard interventions. Similarly, Buchholz et al. (2023) focus on identifying causal models with linear causal mechanisms and nonlinear mixing functions, also utilizing hard interventions. In a more general setting with non-parametric causal mechanisms and mixing functions, von Kügelgen et al. (2023) examine the identifiability of causal models, utilizing multi-environment data from unknown interventions. Similarly, Bagi et al. (2023) explore identifiability of causal models using multi-environment data from unknown interventions. Varici et al. (2023) investigate the identifiability of causal models with nonlinear causal mechanisms and linear mixing functions, considering both hard and soft interventions.
Recent papers have extended explicit hard interventions to soft interventions. Zhang et al. (2023b), the authors indentify the causal model from soft interventions, using sparsity of the adjacency matrix as an inductive bias. For implicit models, however, soft interventions give rise to new challenges. Identifiability is more difficult with soft interventions, since the effect of causal variables on the observed variables is not clear. Effects could be due either intervention or influences from parent variables on the causal variables. Also note that, as we are retaining implicit modeling, hence, there is no knowledge of parents. In Brehmer et al. (2022), identifiability is theoretically proven for hard interventions. However, practical results for complex causal models with over 10 variables indicate increased ambiguity and confounding factors, making model identification less straightforward. In subsequent sections, we’ll explore limitations of Brehmer et al. (2022)’s identifiability theory with soft interventions and propose using the causal mechanisms switch technique from Pearl (2005); Schölkopf (2019) to help the model identify true causal relationships.
3 Proposed Method
3.1 Structural Causal Models for Intervention Data
A causal model is used to understand and describe the relationships between different variables and how they influence each other through causal mechanisms. A causal model in its most general form can be defined as follows.

### v2

# Implicit Causal Representation Learning via Switchable Mechanisms

@@ Abstract
Learning causal representations from observational and interventional data in the absence of known ground-truth graph structures necessitates implicit latent causal representation learning. Implicit learning of causal mechanisms typically involves two categories of interventional data: hard and soft interventions. In real-world scenarios, soft interventions are often more realistic than hard interventions, as the latter require fully controlled environments. Unlike hard interventions, which directly force changes in a causal variable, soft interventions exert influence indirectly by affecting the causal mechanism. However, the subtlety of soft interventions impose several challenges for learning causal models. One challenge is that soft intervention’s effects are ambiguous, since parental relations remain intact. In this paper, we tackle the challenges of learning causal models using soft interventions while retaining implicit modeling. Our approach models the effects of soft interventions by employing a causal mechanism switch variable designed to toggle between different causal mechanisms. In our experiments, we consistently observe improved learning of identifiable, causal representations, compared to baseline approaches.
@@ 1 Introduction
One of the long-standing challenges in causal representation learning is how to recover the ground-truth causal graph of a system solely from observations. Termed the identifiability of causal models problem, this endeavor is crucial. Without achieving identifiability, we risk erroneously attributing causal relationships to learned representations. Furthermore, statistical models can masquerade as Directed Acyclic Graphs (DAGs) where edges lack causal significance, further complicating our pursuit.
When considering the challenge of identifying causal models, it is known that the Markov condition in graphs is insufficient for this task [26]. Thus, without additional assumptions or data, we find ourselves limited to learning only a Markov Equivalence Class (MEC) of the causal model. Existing works have made different assumptions about availability of ground-truth causal variables labels [34], model parameters [1], availability of paired interventional data [3, 31], and availability of intervention targets [17] to ensure identifiability of causal models.
Figure 1: Difference between hard interventions and soft interventions: As seen in the middle row, hard interventions sever connections with parents. Therefore, an object’s class cannot have any effect on the object’s color when we intervene on color. On the other hand, soft interventions, as shown in the bottom row, allow for such effects.
Interventional data are usually obtained through soft or hard interventions. Hard interventions usually involve controlled experiments and they severe the connection of an intervened variable with its parents [24]. In terms of Structural Causal Models (SCM), hard interventions set the causal mechanism relating a causal variable to its parents, to a constant. Due to ethical or safety reasons, it may not be possible to perform hard interventions in many real-world applications. On the other hand, the effects of soft interventions are more subtle since parent variables can still affect their children. These effects can be modeled by a change in the set of parents, the causal mechanisms, and the exogenous variables [7]. Consequently, hard interventions can also be seen as a special case of soft interventions where the causal mechanism is set to a constant. Illustrated in Figure 1, a prominent challenge in causal representation learning lies in dealing with the ambiguity surrounding the effects of soft interventions. The observed alterations in object colors fail to distinctly elucidate whether they stem from parental influences or the applied interventions.
@@
Additionally, a lack of comprehension regarding causal graphs can pose significant challenges in causal representation learning. In certain applications, the causal graph can be constructed using domain knowledge, allowing us to subsequently learn the causal variables [2, 18, 20]. However, this is not universally applicable, necessitating the direct learning of the causal graph itself. In a Variational AutoEncoder (VAE) framework, there are generally two approaches for causal representation learning: Explicit Latent Causal Models (ELCMs) [34, 1, 35, 37, 17, 15] and Implicit Latent Causal Models (ILCMs) [3]. In ELCMs, the latents are the causal variables and the adjacency matrix of the causal graph is parameterized and integrated into the prior of the latents such that the prior of latents is factorized according to the Causal Markov Condition [27]. This approach to causal representation learning is highly susceptible to becoming stuck in local minima as it is hard to learn representations without knowing the graph, and it is hard to learn the graph without knowing the representations. ILCMs [3] were introduced to circumvent this “chicken-and-egg” problem by using solution functions, which can implicitly model edges in the causal graph rather than explicitly modeling the entire adjacency matrix of the causal model. In ILCMs the latents are the exogenous variables and the there is no explicit parameterization for the graph.
@@
In implicit causal representation learning, the task involves recovering the exogenous variables $\mathcal{E}$ from observed variables $\mathcal{X}$ and learning solution functions. In [3], interventions are assumed to be hard, but this is often unrealistic and does not align with real-world problems. In this paper, we propose a novel approach for Implicit Causal Representation Learning via Switchable Mechanisms (ICRL-SM). We will introduce the causal mechanism switch variable as a way of modeling the effect of soft interventions and identifying the causal variables. Our experiments on both synthetic and large real-world datasets, highlight the efficacy of proposed method in identifying causal variables and promising future directions in implicit causal representation learning. Our key contributions can be summarized as follows:
@@
I. A novel approach for implicit causal representation learning with soft interventions.
@@
II. Employing causal mechanisms switch variable to model the effect of soft interventions.
III. Theory for identifiability up to reparameterization from soft interventions.
@@ 2 Related Work
Causal representation learning has recently garnered significant attention [27, 14]. The primary challenge in this problem lies in achieving identifiability beyond the Markov equivalence class [26]. Solely relying on observational data necessitates additional assumptions regarding causal mechanisms, decoders, latent structure, and the availability of interventional data [22, 28, 36, 25, 15, 1, 40, 13, 34]. Recent works have focused on identifying causal models from collected interventional data instead of making strong assumptions about functions of the causal model. Interventional data facilitates identifiability based on relatively weak assumptions [1, 6, 3, 39, 33]. This type of data can be further categorized based on whether it involves soft or hard interventions, and whether the manipulated variables are observed and specified or latent. Our focus in this paper is on examining soft interventions, encompassing both observed and unobserved variables.
@@
CausalDiscrepancy [38] Nonlinear Full row rank polynomial Soft Explicit Permutation and Affine
CauCA [33] Nonlinear Diffeomorphism Soft Explicit Different based on assumptions
Linear-CD [29] Linear Linear Hard Explicit Permutation
Scale-I [30] Nonlinear Linear Hard/Soft Explicit Scale/Mixed
ILCM [3] Nonlinear Diffeomorphism Hard Implicit Permutation and reparameterization
dVAE [21] Nonlinear Diffeomorphism Hard Implicit Permutation and reparameterization
ICRL-SM (ours) Nonlinear Diffeomorphism Soft Implicit Reparameterization
@@
Table 1 presents a comparison of the assumptions and identifiability results between our proposed theory and other related works on causal representation learning with interventions. In causal representation learning with interventions, one approach assumes a given causal graph and concentrates on identifying causal mechanisms and mixing functions. For instance, Causal Component Analysis (CauCA) [33] explores soft interventions with a known graph. Alternatively, when the graph is not provided, explicit models seek to reconstruct it from interventional data [6, 17], potentially resulting in a chicken-and-egg problem in causal representation learning [3]. Current methods face the challenge of simultaneously learning the causal graph and other network parameters, especially in the absence of information about causal variables or the graph. Addressing these challenges, [3] recently introduced ILCM, which performs implicit causal representation learning exclusively using hard intervention data. In contrast, our approach introduces a novel method for learning an implicit model from soft interventions. [3] describes methods for extracting a causal graph from a learned implicit model, which could be applied to our method as well. In our experiments, we will compare our method with ILCM and dVAE [21], given their implicit nature and similar experimental settings and assumptions. Additionally, to showcase the superiority of our method over explicit models, we will employ explicit causal model discovery methods like ENCO [16] and DDS [5], in conjunction with various variants of $\beta$-VAE.
@@ 2.2 Hard interventions vs Soft interventions
The identification of explicit causal models from hard interventions has been extensively explored. [29] investigate causal disentanglement in linear causal models with linear mixing functions under hard interventions. Similarly, [4] focus on identifying causal models with linear causal mechanisms and nonlinear mixing functions, also utilizing hard interventions. In a more general setting with non-parametric causal mechanisms and mixing functions, [32] examine the identifiability of causal models, utilizing multi-environment data from unknown interventions. Similarly, [2] explore identifiability of causal models using multi-environment data from unknown interventions. [30] investigate the identifiability of causal models with nonlinear causal mechanisms and linear mixing functions, considering both hard and soft interventions.
Recent work has expanded the concept of explicit hard interventions to include soft interventions. In their study, [38] address the identification of causal models from soft interventions, leveraging the sparsity of the adjacency matrix as an inductive bias. However, when dealing with implicit models, soft interventions introduce new complexities. Identifiability becomes more challenging, as the causal effect of variables on observed variables is less apparent. This ambiguity arises from the dual possibility of effects originating from interventions or influences from parent variables on the causal variables. Moreover, in scenarios where implicit modeling is retained, the absence of knowledge about parent variables further complicates identifiability. While [3] theoretically establishes identifiability for hard interventions, practical experiments involving complex causal models with over 10 variables reveal increased ambiguity and confounding factors. Consequently, model identification becomes less straightforward.
3 Methodology
3.1 Data Generating Process
A structural causal model (Definition A1.1) is used to understand and describe the relationships between different variables and how they influence each other through causal mechanisms. A decoder function, $g(\mathbf{z})=\mathbf{x}$, maps a vector of causal values $z$ to observed values $x$. The causal variables $\mathcal{Z}$ are unobserved and the goal is to infer them from interventional data. For each causal variable, a diffeomorphic solution function, $s_{i}:\mathcal{E}_{i}\rightarrow\mathcal{Z}_{i}$, deterministically maps a value for exogenous variable $\mathcal{E}_{i}$ to a value for causal variable $\mathcal{Z}_{i}$. In implicit modeling, we learn the solution functions $s_{i}$ directly, rather than defining them through local mechanisms $f_{i}$. We write $\mathcal{S}$ for the set of all solution functions $s_{i}\in\mathcal{S}$, so $\mathcal{S}:\mathcal{E}\rightarrow\mathcal{Z}.$
Identifying causal models from data can be complex and is often studied within classes of models such as those identifiable up to affine transformations. For example, in the context of nonlinear Independent Component Analysis (ICA), the generative process also involves a mixture function $g$ of latent causal variables $\mathcal{Z}\in\mathbb{R}^{n}$, resulting in observations $\mathcal{X}\in\mathbb{R}^{n}$ [15, 41]. However, a significant distinction between causal representation learning and nonlinear-ICA is that in the former, the causal variables $\mathcal{Z}$ may have complex dependencies. Our objective in this paper is to recover $\mathcal{E}$ from $\mathcal{X}$ and eventually map $\mathcal{E}$ to $\mathcal{Z}$ using solution functions.
Identifying a causal model from observational data is not trivial and requires assumptions on the parameters of the model [1]. Adding information about interventions in addition to observations, helps to identify causal variables by exhibiting the effect of changing a causal variable on the observed variables. An interventional data point $(x,\tilde{x},i)$ includes the pre-intervention observation $x$, the post-intervention observation $\tilde{x}$, and intervention target $i\in\mathcal{I}$ where $\mathcal{I}$ is the set of intervention targets selected from the causal variables. The post-intervention data $\tilde{x}$ is generated by a soft intervention that targets one of the causal variables in $\mathcal{Z}$. To achieve identifiability up to reparametrization, we rely on a series of assumptions within the data generation process, outlined as follows:
Assumption 3.1.
(Data generating assumptions)
1. Atomic Interventions: For every sample $(x,\tilde{x},i)$, only one causal variable is targeted by an intervention.
2. Known Targets: Targets of soft interventions are known.
3. Post-intervention Exogenous Variables: The exogenous variables’ values change only for the corresponding intervened causal variable, while the others maintain their pre-intervention values, thus $e_{i}\neq\tilde{e}_{i}$ if $i\in\mathcal{I}$ ,and $e_{i}=\tilde{e}_{i}$ otherwise.
4. Sufficient Variability: Soft interventions alter causal mechanisms to introduce sufficient variability [15]. These interventions should modify causal mechanisms to ensure non-overlapping conditional distributions of causal variables (refer to Figure A1).
5. Diffeomorphic decoder and causal mechanisms: Diffeomorphism guarantees no information loss and avoids abrupt changes in the function’s image.
The known targets assumption can be relaxed in applications where such data is not available and the same procedure in [3] can be used to infer the intervention targets. In fact, in our real-world experiments, intervention targets are not available and based on the nature of the datasets, we hypothesize our causal variables to be object attributes and actions to be intervention targets.
3.2 Causal Mechanisms Switch Variable
The major difference of soft intervention with hard intervention is that post-intervention causal variable $\tilde{\mathcal{Z}_{i}}$ is no longer disconnected from its parents and its causal mechanism $\tilde{s}_{i}$ is affected by the intervention. This is why identifying the causal mechanisms is more difficult for soft interventions. Soft intervention data yield fewer constraints on the causal graph structure than hard intervention data. For more details refer to string diagrams of soft and hard interventions depicted in Figure A5. Figure 2(b) shows our main generative model. It includes a data augmentation step that adds the intervention displacement $\tilde{x}-x$ as an observed feature that directly represents the effect of a soft intervention in observation space.
Augmented implicit causal model
To model the effect of soft interventions, we introduce the causal mechanism switch variable $\mathcal{V}$ [26]. By leveraging $\mathcal{V}$, we can effectively switch to the pre-intervention causal mechanisms within post-intervention data. This facilitates the model’s ability to solely focus on discerning alterations in the intrinsic characteristics of each causal variable. These changes are encapsulated within their respective exogenous variables, aiding the model in learning the causal relationships more accurately. We propose to use a modulated form of $\mathcal{V}$ to model the soft intervention effects on each causal variable as an additive effect with a nonlinear function $h_{i}$ such that $\forall i,\>\tilde{\mathcal{Z}_{i}}=\tilde{s}_{i}(\tilde{\mathcal{E}}_{i};\tilde{\mathcal{E}}_{/i})=s_{i}(\tilde{\mathcal{E}}_{i};\mathcal{E}_{/i},h_{i}(\mathcal{V}))$. As the parental set for each causal variable is not known, we have to use a modulated form of $\mathcal{V}$ in every causal variable’s solution function and the inclusion of $h_{i}(\mathcal{V})$ enables the model to encompass variations in the parental sets of all causal variables in $\mathcal{V}$. Therefore, there is a switch variable $\mathcal{V}_{i}$ for each causal variable $\mathcal{Z}_{i}$. Adding switch variables to solution functions leads to the concept of an augmented implicit causal model.